# Addressing Diverging Training Costs using Local Restoration for Precise Bird’s Eye View Map Construction

### v3

@@ Abstract
Recent advancements in Bird’s Eye View (BEV) fusion for map construction have demonstrated remarkable mapping of urban environments. However, their deep and bulky architecture incurs substantial amounts of backpropagation memory and computing latency. Consequently, the problem poses an unavoidable bottleneck in constructing high-resolution (HR) BEV maps, as their large-sized features cause significant increases in costs including GPU memory consumption and computing latency, named diverging training costs issue. Affected by the problem, most existing methods adopt low-resolution (LR) BEV and struggle to estimate the precise locations of urban scene components like road lanes, and sidewalks. As the imprecision leads to risky self-driving, the diverging training costs issue has to be resolved. In this paper, we address the issue with our novel Trumpet Neural Network (TNN) mechanism. The framework utilizes LR BEV space and outputs an up-sampled semantic BEV map to create a memory-efficient pipeline. To this end, we introduce Local Restoration of BEV representation. Specifically, the up-sampled BEV representation has severely aliased, blocky signals, and thick semantic labels. Our proposed Local Restoration restores the signals and thins (or narrows down) the width of the labels. Our extensive experiments show that the TNN mechanism provides a plug-and-play memory-efficient pipeline, thereby enabling the effective estimation of real-sized (or precise) semantic labels for BEV map construction. Our codes will be publicly released.
@@ I INTRODUCTION
Bird’s Eye View (BEV) provides an egocentric agent with a comprehensive view of large scenes, facilitating the understanding of a global scene. Owing to its capabilities, it is widely used in constructing urban maps for autonomous driving. Recent suggestions of BEV representation to evaluate latent BEV spaces often employ pillar representation with infinite-height voxels. This approach enables the sensors to share their estimated 3D geometries and promotes collaborative sensing. For example, cameras effectively complement the sparse data distribution of radar and LiDAR with their dense sensing, while LiDAR robustly and broadly collects surround geometry, and radar can estimate accurate poses of objects. Thus, their collaborative fusion in a high-dimensional BEV space enhances a robotic system’s scene understanding, as evidenced in Kim et al.’s study [10].
However, existing BEV fusions face a significant hurdle in inferring high-resolution (HR) BEV representations (See Fig. 1 for various BEV resolutions). The heavy architecture of the frameworks, which includes multiple encoders, decoders, feature pyramid networks [14], and a feature lifting [21] module, cause substantial memory consumption and computational bottlenecks. Consequently, when a BEV is configured with a higher resolution or heavy mechanism like attention [30], both backpropagation memory and computing latency increase significantly. Because the issue, which we term the diverging training costs, limits the learning of HR BEVs with high costs, recent self-driving agents rely on low-resolution (LR) BEVs exposing them to risky behaviors such as crossing stop lines or road lanes. For example, in Fig. 1, despite depicting the same section of the road, the low-resolution maps indicate that the vehicle remains within its lane, while the high-resolution maps definitively show that the vehicle crosses the lane boundary. Since this misunderstanding endangers safe motion planning, we have to address this problem. To this end, it is necessary to explore efficient pipelines for representing a HR BEV, thereby it becomes generally available in self-driving fields.
@@
[…]
Figure 1: Local Restoration for Precise Map Construction. “m/px” means meter per pixel. Precise map construction is crucial for safe autonomous driving. However, its training is significantly limited by huge memory due to deep and bulky BEV architecture. Our plug-and-play mechanism resolves the issue, allowing for efficient and precise map construction.
Figure 2: Overview of Our Proposed Methods. Our trumpet neural network take in LiDAR ($\mathbf{z}^{LR}_{p}$) and Camera ($\mathbf{z}^{LR}_{i}$) BEV features. After the cross-modal fusion neck ($\mathbf{B}_{\psi}$) fuse and enhance them, our Local Restoration ($S$) up-samples and restores a unified HR BEV feature. Then decoding CNNs ($\mathbf{D}_{\phi}$) estimate an HR semantic map.
In this paper, we introduce a plug-and-play, memory-efficient approach, Trumpet Neural Network (TNN). Our method evaluates camera and LiDAR BEV features in a sparsely voxelized 3D space. Then, after feature enhancement, TNN up-samples and restores local details of BEV as in Fig. 3. This pipeline aims to construct HR BEV space, akin to the amplified buzzing sound of a trumpet. Our approach alleviates diverging training costs while achieving performance gains, thereby establishing a new baseline of BEV representation. With the improvements, the proposed mechanism provides an effective model training, and the achievement facilitates the learning of HR map construction. In our experiments, we focus on examining the compatibility of TNN with BEV representation using straightforward vanilla frameworks. Our contributions are summarized as:
  • Our proposed mechanism addresses the issue of diverging training costs, thereby facilitating efficient training of models for representing high-resolution BEVs, which are crucial for safe self-driving.
  • TNN is the first framework to utilize BEV restoration, which up-samples small-sized BEV features, thins the semantic labels, and restores their aliasing and blocky artifacts. Owing to this novel approach, TNN constructs maps in LR BEV and decodes them to precise HR BEV, effectively addressing the diverging training costs.
  • We extensively investigate the compatibility of the TNN mechanism with LiDAR, camera, and their fusion.
@@ II RELATED WORKS
The learning methodology for the fusion of LiDAR and camera has shown promising demonstrations in map construction [24, 8, 9, 25, 36, 13, 34, 12, 20, 1, 33, 18, 31, 35, 7, 32, 33, 10, 26, 11, 37, 4]. Borse et al. suggested X-Align [1] and demonstrates that domain alignment enhances the robustness of BEV features against noisy and blurred effects. Kim et al. [10] introduced a methodology for broad BEV perception. Wang et al. [32] suggested an effective transformer architecture for multi-modal fusion. Kim et al. suggested CRN [11], a camera-radar fusion utilizing prospectively projected camera images for enriching semantic camera BEV features. Chang et al. proposed a BEVMap [4], which utilizes a pre-built map to address the inaccurate placement of camera BEV features in 3D space. HDMapNet [12] suggests a general framework for HD map construction with BEV segmentation. LiDAR2Map [33] proposed a distillation strategy to construct HD maps with LiDAR-only modality. Shin et al. suggested InstaGraM [29], a real-time HD map constructor. Qiao et al. proposed BeMapNet [22], a map constructor with bezier curves, and Machmap [23], an end-to-end solution for BEV map construction.
@@
Considering the diverging training costs of HR BEV fusion systems, there is no alternative but to use the task-specific BEV scopes as the dense and broad urban mapping requires large-sized tensor templates of BEV features. For example, if a map constructor uses a broad and dense (-50m, 50m, 0.1m/px) scope, it would encounter severe computational overheads and memory consumption due to the size of BEV features ($1000\times 1000\times C$). Therefore, utilizing the HR variables with limited hardware resources necessitates narrowing perception ranges to mitigate diverging training costs. To address the problem, we suggest Trumpet Neural Network, which processes coarsely voxelized BEVs and outputs an up-sampled HR BEV representation.
Figure 3: Comparison of TNN Restoration to conventional methods. We use a (-50m, 50m) BEV scope with 2.0m/px LR BEV and 0.5m/px HR BEV resolutions. The hand-crafted Nearest, Bilinear, and Bicubic methods suffer from the incomplete restoration of aliasing and blocky artifacts, compared to the TNN, which utilizes learnable restoration.
@@ III PROBLEM FORMULATION
This section outlines the formulation of TNN ($\mathbf{G}_{\theta}$), where $\theta$ denotes parameter. First, we revisit the conventional representation of a fused BEV feature map. Then, we describe the TNN mechanism and clarify the challenges of its implementation. Our formulations are under the assumption that a camera ($\mathbf{z}_{i}\in\mathbb{R}^{C_{i}}$), a LiDAR ($\mathbf{z}_{p}\in\mathbb{R}^{C_{p}}$), and a fused ($\mathbf{z}\in\mathbb{R}^{C_{f}}$) BEV feature are prepared from $K$ images of multiple cameras ($\mathbf{I}$) and $N$ points of LiDAR ($\mathbf{P}$) as in Liu et al’s method [18]. We respectively notate $D\times W$ and $d\times w$ as HR and LR sizes.
Table I: Comparisons of up-sampling methods and computational efficiency using ×$\times 4$ models. The Learnable Deconvolution and Pixel Shuffle show effective restoration. Deconvolution is an over-determined form of Pixel Shuffle that consumes large computational costs without improving performance, which is why we utilize Pixel Shuffle.
  Method Latency (ms) GPU Mem. (GB) mIoU
  Nearest 151 3.6 62.9
Bilinear 155 3.6 63.5
Bicubic 158 3.6 63.6
Deconvolution 167 3.9 70.9
Pixel Shuffle 164 3.8 70.9
Table II: Plug-and-play performance gains of TNN. We plug ×$\times 4$ TNN into the camera, LiDAR, and LiDAR-camera fusion modalities, employing three distinct backbones to investigate TNN’s compatibility.
  Method Modality ×$\times 4$ TNN mIoU
  Swin-T [17] + LSS [21] Camera 57.1
61.4
  Sparse Conv. [16] LiDAR 48.6
68.7
  BEVFusion [18] Lidar-camera Fusion 62.7
70.9
@@ III-A Convention of BEV Fusion
Figure 4: Restoration Comparisons on varying up-sampling factors. We use a (-50m, 50m) BEV scope with 0.5m/px, 1.0m/px, 2.0m/px, 4.0m/px LR BEV resolutions for ×1,×2,×4$\times$ 1$\times$ 2$\times$ 4× 1 , × 2 , × 4, and ×$\times 8$ TNN, respectively. Their HR BEV space utilizes 0.5m/px HR resolution. The ×$\times 1$ and ×$\times 2$ TNNs fail to learn high-level scene features (Black ROIs). In contrast, ×$\times 8$ TNN suffers from inaccurate prediction (Green ROIs) caused by severe loss of prior, compared to the ×$\times 4$ TNN.
@@ III-B Our Strategy
A semantic BEV map for driving perception is composed of high-level global features. Thus, learning restoration of the aliased and blocky BEV can fulfill the requirements for large-scale perception in urban environments thanks to its patterned and simple-to-estimate shapes [19]. To learn local restoration, our proposed mechanism leverages Pixel Shuffle [27] as:
$\hat{\mathbf{z}}=S(\mathbf{B}_{\psi}(f_{\nu}(\mathbf{z}^{LR}_{p},\;\mathbf{z}^{LR}_{i}));\varphi),\;\;S=\mathcal{PS}\circ f_{\varphi},$ (3)
$S(\cdot;\varphi):\mathbb{R}^{d\times w}\mapsto\mathbb{R}^{D\times W}$ is a sequential operation of neural networks ($f_{\varphi}$) followed by Pixel Shuffle ($\mathcal{PS}$), $\mathbf{z}^{LR}\in\mathbb{R}^{C_{f}}$ is an LR BEV feature. With our HR BEV evaluation in Eq. 3, $f_{\varphi}$ learns the restored HR BEV representation, and the composed neural nets learn compressed semantic labels as in the zoom-in regions of Fig. 7. Note that the BEV of urban environments has categorized landscapes as Lynch et al.’s analysis [19]. LR BEV’s representation power for the approximation is enough. Our suggested mechanism answers semantic labels in LR BEV space and up-samples them to HR BEV space by restoring aliased and blocky features.
@@ IV METHOD
The overall pipeline is illustrated in Fig. 2. Our framework takes in $K$ images and $N$ LiDAR points to estimate a semantic BEV map. It adheres to the conventions established by BEVFusion [18] for extracting camera and LiDAR BEV features. Subsequently, the TNN unifies the two BEV features into a unified BEV feature map and enhances it with a fusion neck. Afterward, it deploys local restoration to construct HR BEV and decode them into a semantic map.
@@ IV-A Evaluation of Low-resolution BEV
Table III: Analysis of TNN’s effectiveness tendency with increasing up-scaling factors. Compact, smaller LR BEV space (×4,×8$\times$ 4$\times$ 8× 4 , × 8) effectively approximate urban scenes. However, excessively small-sized BEV spaces often exhibit severely aliased and blocky features as indicated by the mIoU gap between ×$\times 4$ and ×$\times 8$ scales.
  Scale Factor mIoU Latency GPU Mem.
(ms) (GB)
  Baseline 62.7 176 3.1
×$\times 1$ 66.3 218 3.8
×$\times 2$ 69.6 187 4.2
×$\times 4$ 70.9 164 3.8
×$\times 8$ 70.7 153 3.7
Table IV: Comparison of LR to HR BEV segmentations. “mIoU ($r$)” denotes the evaluated mIoU of “$r$” BEV resolution. “TNN” method is pre-trained on 2.0m/px, and fine-tuned on 0.5m/px BEV segmentation.
  Method Trained mIoU mIoU
BEV Resolution (2.0m/px) (0.5m/px)
  LR BEV 2.0m/px 70.8 -
HR BEV 0.5m/px - 62.7
×$\times 4$ TNN 2.0m/px→ 0.5m/px 70.8 70.9
IV-B Trumpet Neural Network
TNN simultaneously carries out up-sampling and the estimation of local details in the BEV representation ($\hat{\mathbf{z}}^{HR}$) with our proposed operator ($S(\cdot;\varphi,s)$) as:
@@ IV-C Manipulation of BEV Scope and Resolution
Spatial Pooling. Pooling with kernels, such as CNN, and max pooling enlarges the size of the receptive field and the coverage of BEV ranges. Given that $\mathbf{U}=[u,v]$ represents a pixel coordinate, $\chi=\{[i,j]\;|\;i\in\chi_{i},\;j\in\chi_{j}\}$ is a set of relative coordinates for a pooling kernel with $|\chi_{i}|\times|\chi_{j}|$ size, and $\hat{\mathbf{z}}$ is the pooled feature map, the BEV space of a feature fiber $\hat{\mathbf{z}}(\mathbf{U;\chi})$ with the pre-defined $\mathbf{R}$ scope in pixel and $\mathbf{r}=[r_{i},r_{j}]$ resolution are re-defined $(\cdot)^{\prime}$ as:
$\mathbf{R}^{\prime}=\mathbf{R}_{u}\times\mathbf{R}_{v},\;\;\mathbf{r}^{\prime}=[\;\frac{r_{i}}{|\chi_{i}|},\;\frac{r_{j}}{|\chi_{j}|}\;],$ (7)
$\textrm{where}\;\mathbf{R}_{u}=\{u+i\;|\;i\in\chi_{i}\},\;\mathbf{R}_{v}=\{v+j\;|\;j\in\chi_{j}\}.$ (8)
As in Eq. 7, processing BEV features with max-pooling manipulates the pixel-wise coverage and size of the receptive field. Therefore, the BEV feature’s resolution is down-sampled while becoming a global representation.
@@
Local Restoration. The operation consists of pixel shuffle ($\mathcal{PS}$) and CNNs ($f_{\varphi}$). In analytical point of view, $f_{\varphi}(\cdot):\mathbb{R}^{h\times w\times C}\mapsto\mathbb{R}^{h\times w\times s^{2}C}$ projects latent geometry features with $s$-scaled scope in their channels, $\mathcal{PS}(\cdot):\mathbb{R}^{sh\times sw\times C}$ reshapes them to an HR BEV. Thus, local restoration preserves BEV scope ($\mathbf{R}$) but manipulates resolution ($r$). It estimates interpolated BEV features from LR BEVs restoring local details, due to the optimization policy that imposes the model to align with the HR semantic BEV maps’ ground truth.

### v4

# Addressing Diverging Training Costs using BEVRestore for High-resolution Bird’s Eye View Map Construction

@@ Abstract
Recent advancements in Bird’s Eye View (BEV) fusion for map construction have demonstrated remarkable mapping of urban environments. However, their deep and bulky architecture incurs substantial amounts of backpropagation memory and computing latency. Consequently, the problem poses an unavoidable bottleneck in constructing high-resolution (HR) BEV maps, as their large-sized features cause significant increases in costs including GPU memory consumption and computing latency, named diverging training costs issue. Affected by the problem, most existing methods adopt low-resolution (LR) BEV and struggle to estimate the precise locations of urban scene components like road lanes, and sidewalks. As the imprecision leads to risky motion planning like collision avoidance, the diverging training costs issue has to be resolved. In this paper, we address the issue with our novel BEVRestore mechanism. Specifically, our proposed model encodes the features of each sensor to LR BEV space and restores them to HR space to establish a memory-efficient map constructor. To this end, we introduce the BEV restoration strategy, which restores aliasing, and blocky artifacts of the up-scaled BEV features, and narrows down the width of the labels. Our extensive experiments show that the proposed mechanism provides a plug-and-play, memory-efficient pipeline, enabling an HR map construction with a broad BEV scope. Our code will be publicly released.
@@ I INTRODUCTION
Bird’s Eye View (BEV) provides an egocentric agent with a comprehensive view of large scenes, facilitating the understanding of a global scene. Owing to its capabilities, it is widely used in constructing urban maps for autonomous driving. Recent BEV representation of deep nets applies a memory-efficient 2D space with infinite-height voxels, and associates a unified space for various modalities, such as radar, LiDAR, and camera. This approach enables the sensors to share their estimated 3D geometries and promotes collaborative sensing. For example, cameras effectively complement the sparse data distribution of radar and LiDAR with their dense sensing, while LiDAR robustly and broadly collects surround geometry, and radar can estimate accurate poses of objects. Thus, their collaborative fusion in a high-dimensional BEV space enhances a robotic system’s scene understanding [10].
@@
(a) High-resolution map construction using BEVRestore.
(b) Training memory efficiency.
(c) BEV Seg. performance.
Figure 1: “m/px” means meter per pixel. Training with HR BEV features consumes a huge memory size due to the bulky architecture. Our plug-and-play BEVRestore (Fig. 1(a)) addresses the issue, allowing for costly efficient map construction (Fig. 1(b)) and enhancing feature encoding (Fig. 1(c)).
However, existing BEV fusions face a significant hurdle in inferring high-resolution (HR) BEV representations (See Fig. 1(a) for low and high BEV resolutions). Their heavy architecture, which includes multiple encoders, decoders, feature pyramid networks [15], and a feature lifting [22] module, cause substantial memory consumption and computational bottlenecks. Consequently, when a BEV is configured with a higher resolution or heavy mechanism like attention [30], both backpropagation memory and computing latency increase significantly. Because the issue, which we term the diverging training costs, limits the learning of HR BEVs with high costs, recent self-driving agents rely on low-resolution (LR) BEVs exposing them to risky behaviors such as crossing stop lines or road lanes. For example, in Fig. 1, despite depicting the same section of the road, the low-resolution maps indicate that the vehicle remains within its lane, while the high-resolution maps definitively show that the vehicle crosses the lane boundary. Since this misunderstanding endangers safe motion planning, we have to address this problem. To this end, it is necessary to explore efficient pipelines for representing an HR BEV, thereby it becomes generally available in self-driving fields.
In this paper, we introduce a plug-and-play, memory-efficient approach, BEVRestore. Our method encodes the context features of each sensor into an LR BEV space. Then, after feature enhancement, BEVRestore reconstructs local details. Our approach alleviates diverging training costs while achieving performance gains establishing a new baseline for BEV map construction. In our experiments, we focus on examining the performance improvement and compatibility of BEVRestore across sensors, BEV encoders, and map construction tasks including BEV segmentation and HD map construction. Our contributions are summarized as:
  • Our proposed mechanism addresses diverging training costs, thereby facilitating efficient training of models for representing high-resolution BEVs, which are crucial for safe self-driving.
  • BEVRestore is the first framework to utilize restoration of low-resolution BEV, which up-samples BEV features, thins the semantic labels, and restores their aliasing and blocky artifacts. Owing to this costly novel approach, BEVRestore efficiently constructs HR BEV maps addressing the diverging training costs.
  • Our approach improves map construction performance. We extensively investigate the property in addition to its compatibility across various sensors, BEV encoders, and map construction tasks.
@@ II RELATED WORKS
Fusion of LiDAR-camera BEV has shown promising demonstrations in map construction [25, 8, 9, 36, 14, 34, 13, 21, 1, 33, 19, 31, 35, 7, 32, 33, 10, 26, 11, 37, 4]. Borse et al. suggested X-Align [1] and showed that domain alignment enhances the robustness of BEV features against noisy and blurred effects. Kim et al. [10] introduced a methodology for broad BEV perception. Wang et al. [32] suggested an effective transformer architecture for multi-modal fusion. Kim et al. suggested CRN [11], a camera-radar fusion utilizing prospectively projected camera images for enriching semantic camera BEV features. Chang et al. proposed a BEVMap [4], which utilizes a pre-built map to address the inaccurate placement of camera BEV features in 3D space. HDMapNet [13] suggests a general framework for HD map construction with BEV segmentation. LiDAR2Map [33] proposed a distillation strategy to construct HD maps with LiDAR-only modality. Shin et al. suggested InstaGraM [29], a real-time HD map constructor. Qiao et al. proposed BeMapNet [23], a map constructor with bezier curves, and Machmap [24], an end-to-end solution for BEV map construction.
@@
Considering the diverging training costs of HR BEV fusion systems, there is no alternative but to use the task-specific BEV scopes as the dense and broad urban mapping requires large-sized tensor templates of BEV features. For instance, if a map constructor uses a broad and dense (-50m, 50m, 0.1m/px) scope, it would encounter severe computational overheads and memory consumption due to the size of BEV features ($1000\times 1000\times C$). Therefore, utilizing the HR variables with limited hardware resources necessitates narrowing perception ranges to mitigate diverging training costs. To address the problem, we suggest BEVRestore, which processes coarsely voxelized BEVs and outputs an up-sampled HR BEV representation.
@@ III PROBLEM FORMULATION
This section outlines the formulation of BEVRestore ($\mathbf{G}_{\theta}$), where $\theta$ denotes parameter. First, we revisit the conventional representation of a fused BEV feature map. Then, we describe the BEVRestore mechanism and clarify the challenges of its implementation. Our formulations are under the assumption that a camera ($\mathbf{z}_{i}\in\mathbb{R}^{C_{i}}$), a LiDAR ($\mathbf{z}_{p}\in\mathbb{R}^{C_{p}}$), and a fused ($\mathbf{z}\in\mathbb{R}^{C_{f}}$) BEV feature are prepared from $K$ images of multiple cameras ($\mathbf{I}$) and $N$ points of LiDAR ($\mathbf{P}$) as in Liu et al’s method [19]. We respectively notate $D\times W$ and $d\times w$ as HR and LR sizes.
@@ III-A Convention of BEV Fusion
Figure 2: Overview of Our Proposed Methods. Our suggested mechanism take in LiDAR ($\mathbf{z}^{LR}_{p}$) and Camera ($\mathbf{z}^{LR}_{i}$) BEV features. After the cross-modal fusion neck ($\mathbf{B}_{\psi}$) fuse and enhance them, our BEVRestore ($S$) up-samples and restores a unified HR BEV feature. Then decoding CNNs ($\mathbf{D}_{\phi}$) estimate an HR semantic map.
@@ III-B Our Strategy
A semantic BEV map for driving perception is composed of high-level global features due to its broad scope in representing urban scenes. Our approach leverages LR space to prioritize the learning of global features over local details and restores the LR BEV to HR space. To this end, our proposed BEVRestore leverages Pixel Shuffle [27] as:
$\hat{\mathbf{z}}=S(\mathbf{B}_{\psi}(f_{\nu}(\mathbf{z}^{LR}_{p},\;\mathbf{z}^{LR}_{i}));\varphi,s),\;\;S=\mathcal{PS}\circ f_{\varphi},$ (3)
$S(\cdot;\varphi):\mathbb{R}^{d\times w}\mapsto\mathbb{R}^{D\times W}$ is a sequential operation of neural networks ($f_{\varphi}$) followed by Pixel Shuffle ($\mathcal{PS}$), $\mathbf{z}^{LR}\in\mathbb{R}^{C_{f}}$ is an LR BEV feature, $s$ is a down-sizing scale factor. With our HR BEV evaluation in Eq. 3, $f_{\varphi}$ learns the restored HR BEV representation, and the composed neural nets learn compressed semantic labels as in the zoom-in regions of Fig. 6. In urban BEV maps, especially for roads and their surroundings, layouts are often categorized into a few cases with minimal variation. Key points like intersections (“nodes”) and the roads connecting them (“edges”) follow standardized patterns [20] that neural networks can easily approximate. LR BEV has sufficient representation power for this. Our approach generates semantic labels in the LR BEV space and up-samples them to HR BEV, restoring any blocky features.
@@ IV METHOD
The overall pipeline is illustrated in Fig. 2. Our framework takes in $K$ images and $N$ LiDAR points to estimate a semantic BEV map. It adheres to the conventions established by BEVFusion [19] for extracting camera and LiDAR BEV features. Subsequently, the BEVRestore unifies the two BEV features into a unified BEV feature map and enhances it with a fusion neck. Afterward, it deploys BEVRestore to construct HR BEV and decode them into a semantic map.
@@ IV-A Evaluation of Low-resolution BEV
IV-B BEV Restoration
Restoration of BEV simultaneously up-samples and estimates local details of HR BEV representation ($\hat{\mathbf{z}}^{HR}$) with our proposed operator ($S(\cdot;\varphi,s)$) as:
@@ IV-C Manipulation of BEV Scope and Resolution
Spatial Pooling. Pooling with kernels, such as a down-sampling CNN layer, max, and average poolings, enlarges the pixel-wise receptive field size and the corresponding coverage of BEV ranges. Given that $\mathbf{u}=[u,v]$ is a pixel coordinate, $\chi=\{[i,j]\;|\;i\in\chi_{u},\;j\in\chi_{v}\}$ is a relative coordinate set for a pooling kernel with $|\chi_{u}|\times|\chi_{v}|$ size, the pre-defined BEV resolution $\mathbf{r}=[r_{i},r_{j}]$ and the coverage of a BEV pixel $\mathcal{R}(\mathbf{u;r})$ are re-defined $(\cdot)^{\prime}$ as:
$\mathcal{R}^{\prime}(\mathbf{u;\;r^{\prime}})=\bigcup_{[i,j]\in\chi}\{\mathcal{R}(\mathbf{u}+[i,j];\;\mathbf{r}^{\prime})\},$ (7)
$\textrm{where}\;\mathbf{r}^{\prime}=[\;r_{i}\cdot|\chi_{u}|,\;r_{j}\cdot|\chi_{v}|\;].$ (8)
As in Eq. 7, the BEV feature and its coverage are respectively down-sampled and enlarged while becoming a global representation.
Table I: Plug-and-play performance gains for BEV segmentation of various modalities. We plug “×$\times 4$” BEVRestore into the camera, LiDAR, and LiDAR-camera fusion modalities, employing three distinct backbones to investigate BEVRestore’s compatibility.
  Method Modality ×$\times 4$ mIoU
  Swin-T [18] + LSS [22] Camera 57.1
61.4
  Sparse Conv. [17] LiDAR 48.6
68.7
  BEVFusion [19] LiDAR-camera Fusion 62.7
70.9
Table II: Plug-and-play performance gains for HD map construction with camera BEV. We plug “×$\times 4$” BEVRestore into HD map construction models.
  Method ×$\times 4$ mAP mIoU Latency Mem.
(ms) (GB)
  HDMapNet [13] 23.0 34.2 38 1.9
32.1 36.4 58 3.5
  BEVFormer [14] - 40.1 125 9.9
- 41.3 104 2.5
@@
BEVRestore. The operation consists of pixel shuffle ($\mathcal{PS}$) and CNNs ($f_{\varphi}$). In analytical point of view, $f_{\varphi}(\cdot):\mathbb{R}^{h\times w\times C}\mapsto\mathbb{R}^{h\times w\times s^{2}C}$ projects latent geometry features with $s$-scaled scope in their channels, $\mathcal{PS}(\cdot):\mathbb{R}^{sh\times sw\times C}$ reshapes them to an HR BEV. Thus, BEVRestore preserves BEV scope ($\mathbf{R}$) but manipulates resolution ($r$). It estimates interpolated BEV features from LR BEVs restoring local details, due to the optimization policy that imposes the model to align with the HR semantic BEV maps’ ground truth.
IV-D Implementation Details
Our implementations are built on top of mmdetections [5], [6]. We use the Swin-T (Tiny) [18] model and feature lifting module [22] to encode the camera BEV. We adopt sparse convolution [17] as a LiDAR backbone. To fuse the LiDAR and the camera BEV features, we use a fully convolutional layer with Feature Pyramid Networks (FPN) [15]. The data augmentations we used are the same methods as the baseline work [19]. Detailed model and training configurations are shown in our publicly available code.
Table III: Comparison of LR to HR BEV segmentations. “mIoU ($r$)” denotes the evaluated mIoU of “$r$” BEV resolution. “BEVRestore” and “BEVRestore-S (Small)” are pre-trained on 2.0m/px, and fine-tuned on 0.5m/px BEV space. To compose “BEVRestore-S”, We downsize the number of parameters of BEVRestore and decoding nets for pure exploration of the effectiveness of LR BEV space
  Method Trained mIoU mIoU Param.
BEV Resolution (2.0m/px) (0.5m/px) (M)
  LR BEV 2.0m/px 70.8 - 46.8
HR BEV 0.5m/px - 62.7 46.8
×$\times 4$ BEVRestore-S 2.0 → 0.5m/px 70.8 64.7 44.5
×$\times 4$ BEVRestore 2.0 → 0.5m/px 70.8 70.9 148.3
IV-E Training
Our optimization aims to implement a modular BEVRestore mechanism. We first train a base model, which utilizes LR BEV. Afterward, we freeze the parameters before BEVRestore and fine-tune BEVRestore and decoder to use HR BEV space. This strategy enables attaching varying-scaled learnable up-sampling to the latter of the model. Furthermore, the designs address the concern that high-resolution BEV consumes large memory space generating partial backpropagation memory.
V EXPERIMENTAL RESULTS
V-A Quantitative Results
Setup. We apply a ×$\times 4$ BEVRestore, (-50m, 50m) BEV scope, 2.0m/px LR and 0.5m/px HR BEV resolutions in Table I, Table III, Table IV, and Table VI. In Table II, we use (-30m, 30m) and (-15m, 15m) BEV scopes with 0.6m/px LR and 0.15m/px HR BEV resolutions. Gray shading in each table represents our base model.
Cost Measurement. We examine the inference time latency and the training time maximum memory based on a single batch using an NVIDIA A100 80GB GPU.
Table IV: Comparisons of up-sampling methods and computational efficiency using ×$\times 4$ models. The Learnable Deconvolution and Pixel Shuffle show effective restoration. Deconvolution is an over-determined form of Pixel Shuffle that consumes large computational costs without improving performance, which is why we utilize Pixel Shuffle.
  Method Latency (ms) Mem. (GB) mIoU
  Nearest 151 3.6 62.9
Bilinear 155 3.6 63.5
Bicubic 158 3.6 63.6
Deconvolution 167 3.9 70.9
Pixel Shuffle 164 3.8 70.9
Table V: Effectiveness tendency with increasing up-scale factors. Smaller LR BEV space is effective in learning the context of huge scenes. However, excessively down-sized BEV space limits the improvement as indicated by the mIoU gap between ×$\times 4$ and ×$\times 8$ scales.
  Method Resolution mIoU Latency Mem. Param.
(ms) (GB) (M)
  Baseline 0.5m/px 62.7 176 3.6 46.8
×$\times 4$ BEVRestore-S 2.0 $\rightarrow$ 0.5m/px 64.7 157 2.7 44.5
×$\times 1$ BEVRestore 0.5 $\rightarrow$ 0.5m/px 66.3 218 5.1 129.4
×$\times 2$ BEVRestore 1.0 $\rightarrow$ 0.5m/px 69.6 187 4.2 138.9
×$\times 4$ BEVRestore 2.0 $\rightarrow$ 0.5m/px 70.9 164 3.8 148.3
×$\times 8$ BEVRestore 4.0 $\rightarrow$ 0.5m/px 70.7 153 3.7 157.7
Compatibility. In Table I, we evaluate the compatibility of the BEVRestore mechanism across various modalities, including LiDAR, camera, and LiDAR-camera fusion. As indicated in the tables, all modalities exhibit performance gains, which can be attributed to the superior representation power of LR BEV for semantic map construction, coupled with BEVRestore to evaluate HR BEV. In Table II, we plug BEVRestore to HDMapNet [13] and a query-based camera BEV encoder BEVFormer [14] to explore BEVRestore’s compatibility for the task of vectorized semantic map construction. We follow the evaluation protocol of HDMapNet [13] measuring mAP under [0.2, 0.5, 1.0] chamfer distance thresholds, and mIoU of directly estimated semantic maps. As shown in the table, our mechanism improves the map construction showing 9.1% mAP and 2.2% mIoU gains for vectorized and rasterized map construction of HDMapNet, 1.2% for rasterized construction of BEVFormer. Furthermore, we find that the memory consumption of BEVFormer is significantly reduced owing to the BEVRestore mechanism’s utilization of LR BEV space. However, for a light model without resource-intensive modules, BEVRestore may cause an overall increase in costs as shown in the HDMapNet’s latency and memory consumption as our mechanism imposes additional costs for restoration. Despite the inevitable amount of overhead, BEVRestore prevents cost divergence achieving performance gains.
BEVRestore. In Table III, we analyze the restoration performance. For this purpose, we establish a baseline (or BEVFusion) trained under two different conditions: a 0.5m/px HR BEV space and a 2.0m/px LR BEV space. Additionally, we include a base model equipped with our proposed method for comparison. BEVRestore and BEVRestore-S (Small) use 2.0m/px for LR BEV and up-samples it to 0.5m/px BEV resolution. We composed BEVRestore-S to explore LR BEV space’s effectiveness under the smaller representation power by downsizing the number of parameters of BEVRestore and decoding nets. As shown in the first and second rows of the table, compared to 0.5m/px resolution, BEV map construction with 2.0m/px LR BEV offers superior perception for understanding huge urban scenes owing to the LR space’s condensed learning of fine details, allowing the model to focus on broader patterns and overall structures [12]. As in the third and fourth rows, Our mechanism leverages this property by locally restoring the up-sampled BEV representation. In Table IV, we compare trainable up-samplings with the hand-crafted approaches: ‘Nearest’ neighbor, ‘Bilinear’, and ‘Bicubic’ interpolations, to validate the importance of the learnable BEVRestore. The comparisons reveal three insights. 1) Learnable up-samplings (Pixel shuffle, deconvolution) show superior BEV representation compared to handcrafted methods (Nearest neighbor, Bilinear, and Bicubic interpolations). This suggests that non-linear aliasing and blocky artifacts in BEV can be effectively restored using neural networks. 2) Deconvolution is an over-determined form of pixel shuffle as analyzed by Shi et al. [28]. Both methods yield similar performance although pixel shuffle is more cost-effective. 3) The BEVRestore mechanism’s learnable up-sampling approach offers a reasonable trade-off between performance and costs. Compared to non-learnable methods, it achieves 14.2% mIoU gain with an additional 0.4GB memory consumption and a 6ms latency increase in comparison to pixel shuffle over Bicubic interpolation.